# LLMs may Dominate Information Access: Neural Retrievers are Biased Towards LLM-Generated Texts

### v2

@@ 4. RQ3: The Cause of Source Bias
In this section, we delve deeper into the cause of the observed source bias. We aim to confirm whether LLM-generated texts, having less noise and more concentrated semantic topics, become easier for neural retrieval models to comprehend. We conduct a series of analyses from the perspective of text compression and provide valuable theoretical insights.
@@
Figure 8. Comparision of the relative singular value of LLM-generated corpus with different temperatures. The singular values are sorted in descending order from left to right.
4.1. Viewpoint from Compression
In this analysis, we explore the cause of source bias through a compression perspective, drawing inspiration from recent studies that suggest LLMs act as lossless compressors (Delétang et al., 2023). We postulate that LLMs consistently focus on the major part of the information while filtering out the noise. This, however, does not hold true for humans, as human minds tend to get distracted by chance occurrences. To conduct this examination, we assess the number and strength of topics in both the human-written corpus and the LLM-generated corpus. We employ Singular Value Decomposition (SVD) (Klema and Laub, 1980) on these two corpora. The dimension of the SVD corresponds to the maximum number of topics, and the singular value associated with each topic represents its strength. High singular values predominantly capture primary topic information, whereas low singular values indicate noise. To be more specific, we utilize the OpenAI embedding model to obtain embedding matrices for each corpus within the SciFact+AIGC dataset. Subsequently, we perform SVD on these matrices. The singular values are arranged in descending order, and their comparison to the human-written corpus is visualized in Figure 7.
As illustrated in Figure 7, LLM-generated texts exhibit larger singular values at the top large singular values, while smaller singular values at the tail small singular values. This observation suggests that LLM-generated texts tend to have more focused semantics with less noise, rendering them more suitable for precise semantic similarity calculations. In contrast, human-written texts often contain a wider range of topics and higher levels of noise, making them harder for neural IR models to understand. As a result, this difference in semantic concentration may contribute to the observed bias in neural IR models.
Figure 9. Comparision of the PPL of the different corpora.
Furthermore, we conduct an SVD analysis on LLM-generated documents with different temperature coefficients, as shown in Figure 8. Evidently, for various temperature coefficients of Llama2, the head of the SVD consistently displays higher singular values compared to human-written text. However, the tails of the SVD corresponding to higher temperature coefficients align more closely with the human-written text, while lower temperature coefficients result in tails with smaller singular values. This aligns with our previous findings in Section 3.5, that is, lower temperature results in more conservative and predictable text, and higher temperature results in more creative and diverse text. Thus, lower temperature coefficients lead to more substantial compression, greater topic concentration, and a more pronounced source bias.
@@ 4.2. Further Analysis from Perplexity
Considering that most modern neural IR models are grounded on PLMs (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2022), such as BERT (Devlin et al., 2019), Roberta (Liu et al., 2019), and T5 (Raffel et al., 2020), we try to analyze the perplexity of PLMs to further support the conclusion above from the viewpoint of compression that LLM-generated texts can be better understood by PLMs. Perplexity is an important metric for evaluating how well a language model can understand a given text (Azzopardi et al., 2003; Wang et al., 2019). For a specific language model (LM) and a document $d=(d_{0},d_{1},\cdots,d_{S})$ , the log perplexity is defined as the exponentiated average negative log-likelihood of each token in the tokenized sequence of $d$ [^4] For simplicity, we denote the log perplexity as PPL.:
@@
where $S$ represents the token length of each text $d$ and $P_{\text{LM}}(d_{s})$ is the predicted likelihood of the $s$-th token conditioned on the context. Lower perplexity suggests more confidence and understanding of LM to text patterns, while higher perplexity implies greater uncertainty in predictions, often arising from complex or unpredictable text patterns.
Using the most widely-used LM, BERT (Devlin et al., 2019), as an example, we employ it to calculate the PPL for various corpora. As BERT is not an autoregressive LM, we follow standard practices (Wang et al., 2021; Wang and Cho, 2019) to calculate the predicted likelihood of each token conditioned on the other tokens, i.e.,
@@
The distribution of perplexity for different corpus within the SciFact+AIGC dataset is shown in Figure 9.
Notably, LLM-generated texts consistently exhibit significantly lower perplexity (Figure 9), implying enhanced comprehensibility and higher confidence from BERT’s perspective. Consequently, neural retrieval models based on PLMs can more effectively comprehend LLM-generated texts. This enhanced understanding allows for a more accurate prediction of relevance, leading to the observed source bias in favor of LLM-generated texts. Moreover, as shown in Figure 9, Llama2 with a lower temperature coefficient (i.e., $T=0.2$) generates text with a lower perplexity than the higher temperature coefficient (i.e., $T=1.0$). This observation indicates that LLM with a smaller temperature is less prone to generate unexpected or unusual words, thereby enabling a more effective understanding and grasp of language patterns by the PLMs. To conclude, these observations underscore that texts generated by LLM are more easily comprehensible when processed by PLMs in neural IR models, culminating in the observed source bias.
4.3. Theoretical Analysis and Insights
@@
In the theorem below, we introduce two assumptions: Semantic Superiority and Conditional Redundancy, to theoretically establish the sufficient conditions under which $\text{PPL}(d^{G},\mathcal{B})\leq\text{PPL}(d^{H},\mathcal{B})$ holds. Semantic Superiority suggests that the perplexity of human-written texts, when evaluated by humans, is lower than when evaluated by BERT. Conditional Redundancy implies that the perplexity of $d^{G}$, given $d^{H}$, is lower than the perplexity of $d^{H}$ when evaluated directly. This is intuitively true when the information added in generating from $d^{H}$ to $d^{G}$ doesn’t exceed the original information in $d^{H}$. Then we have the following theorem:
Theorem 4.1 ().
@@
if LLM aligns more closely with BERT than with human when predicting $d^{G}$ given $d^{H}$, such that for any $s\in[S],$
[…]
@@
[…]
@@
Proof of Theorem 4.1 can be found in Appendix A.3. In  Theorem 4.1, we assume that there exists an upper bound $\epsilon$ on the increase in perplexity when evaluating $d^{G}$directly, compared to evaluating $d^{G}$ conditioned on $d^{H}$. Moreover, the KL divergence is used to compare the distributions of the document $d^{G}$ conditioned on $d^{H}$ according to the LLM, BERT model, and humans. It is worth to emphasize that inequation  (1) is not the assumption on the understanding capabilities of BERT, LLM, and humans. Instead, this inequation assumes that when predicting $d^{G}$ given $d^{H},$ the predictions by LLM are more closely aligned with those of BERT.

### v3

@@ 4. RQ3: The Cause of Source Bias
In this section, we delve deeper into why neural retrieval models exhibit source bias. Our objective is to determine whether the LLM-generated texts, characterized by reduced noise and more concentrated semantic topics, are inherently easier for neural retrieval models to semantically match. We conduct a series of analyses from the perspective of text compression and provide valuable insights.
@@
4.1. Viewpoint from Text Compression
We first explore the cause of source bias from a compression perspective, drawing inspiration from recent studies that suggest LLMs are lossless compressors (Delétang et al., 2023). We hypothesize that LLMs efficiently focus on essential information, minimizing noise during generation, in contrast to human-written texts, which may include more diverse topics and incidental noise. To verify this, we employ Singular Value Decomposition (SVD) (Klema and Laub, 1980) to compare topic concentration and noise in human-written and LLM-generated texts. The dimension of the SVD corresponds to the maximum number of topics, and the singular value associated with each topic represents its strength. High singular values predominantly capture primary topic information, whereas low singular values indicate noise.
Specifically, we utilize the OpenAI embedding model to obtain embedding matrices for each corpus in the SciFact+AIGC dataset and then conduct SVD. The resulting singular values are arranged in descending order, and their comparison to the human-written corpus is visualized in Figure 6. As we can see, LLM-generated texts exhibit larger singular values at the top large singular values, while smaller singular values at the tail small singular values. This observation suggests that LLM-generated texts tend to have more focused semantics with less noise, rendering them more suitable for precise semantic matching. In contrast, human-written texts often contain a wider range of latent topics and higher levels of noise, making them harder for neural retrievers to understand. As a result, this difference in semantic concentration may contribute to the observed source bias in neural retrievers.
Figure 7. Comparision of the PPL of the different corpus.
@@ 4.2. Further Analysis from Perplexity
Considering that most modern neural retrievers are grounded on PLMs (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2023a), such as BERT (Devlin et al., 2019), Roberta (Liu et al., 2019), and T5 (Raffel et al., 2020), we analyze the perplexity of PLMs to further support the conclusion above from the viewpoint of compression that LLM-generated texts can be better understood by PLMs. Perplexity is an important metric for evaluating how well a language model can understand a given text (Azzopardi et al., 2003; Wang et al., 2019). For a specific language model (LM) and a document $d=(d_{0},d_{1},\cdots,d_{S})$, the log perplexity is defined as the exponentiated average negative log-likelihood of each token in the tokenized sequence of $d$ [^5] For simplicity, we denote the log perplexity as PPL.:
@@
where $S$ is the token length of text $d$ and $P_{\text{LM}}(d_{s})$ is the predicted likelihood of the $s$-th token conditioned on the context. Lower perplexity suggests more confidence and understanding of LM for text patterns, while higher perplexity implies greater uncertainty in predictions, often arising from complex or unpredictable text patterns.
Using the most widely-used LM, BERT (Devlin et al., 2019), as an example, we employ it to calculate the PPL for different corpus. As BERT is not an autoregressive LM, we follow standard practices (Wang et al., 2021; Wang and Cho, 2019) to calculate the likelihood of each token conditioned on the other tokens, i.e.,
@@
The distribution of perplexity for different corpus in the SciFact+AIGC dataset is shown in Figure 7. Notably, LLM-generated texts consistently exhibit significantly lower perplexity, indicating enhanced comprehensibility and higher confidence from BERT’s perspective. Consequently, PLMs-based neural retrievers can more effectively model the semantics of LLM-generated texts, leading to the observed source bias in favor of LLM-generated texts.
In Appendix C, we also provide a theoretical analysis to illustrate and verify the observation in Figure 7 that LLM-generated texts have a smaller perplexity than human-written texts.
5. RQ4: Mitigating Source Bias
In this section, we propose a simple but effective approach to mitigate source bias by introducing a debiased constraint to the optimization objective. In this way, we can force the neural IR models to focus on modeling semantic relevance rather than the inherent semantic shortcut of the LLM-generated content.
(a) Results on mixed human-written and Llama2-generated corpora
(b) Results on mixed human-written and ChatGPT-generated corpora
Figure 8. Performance comparison (NDCG@$1$) of neural models on SciFact+AIGC dataset with different debiased co-efficient $\alpha$. The grey dashed line represents Relative $\Delta=0$.
5.1. Our Method: A Debiased Constraint
Our earlier findings of source bias indicate that neural retrievers tend to rank LLM-generated documents in higher positions. Thus, the motivation of our debiased method is straightforward, which is to force the retrieval models to focus on modeling the semantic relevance and not assign higher predicted relevance scores to the LLM-generated documents. Specifically, following the practice in Section 2.2, we first generate the corresponding LLM-generated corpus $\mathcal{C}^{G}$ for the original human-written training corpus $\mathcal{C}^{H}$. In this way, we can get the new paired training data $\mathcal{D}=\{(q_{m},d^{H}_{m},d^{G}_{m})\}_{m=1}^{M}$, where each element $(q_{m},d^{H}_{m},d^{G}_{m})$ is a <query, human-written document, LLM-generated document> triplet. $d^{H}_{m}$ and $d^{G}_{m}$ are the corresponding human-written and LLM-generated relevant documents for the query $q$, respectively. Then we introduce the debiased constraint, which can be defined as
where $\hat{r}(q,d^{G};\Theta)$ and $\hat{r}(q,d^{H};\Theta)$ are the predicted relevance scores of $(q,d^{G})$ and $(q,d^{H})$ by the retrieval models with parameters $\Theta$, respectively. This constraint can penalize biased samples when the predicted relevance score of $(q,d^{G})$ is greater than that of $(q,d^{H})$.
Based on the debiased constraint defined in (1), we can define the final loss for training an unbiased neural retriever:
where the $\mathcal{L}_{\text{rank}}$ can be any common-used loss for the ranking task, e.g., contrastive loss or regression loss (Zhao et al., 2023a; Guo et al., 2020, 2022). And $\alpha$ is the debiased co-efficient that can balance the ranking performance and the degree of the source bias. The larger $\alpha$ indicates the greater penalty on the biased samples, leading to the retriever being more likely to rank the human-written texts in higher positions.
5.2. Results and Analysis
To evaluate the effectiveness of our proposed debiased method, we equip the debiased constraint defined in Eq. (1) to two representative neural retrievers: ANCE (Xiong et al., 2020) and BERM (Xu et al., 2023). In the experiments, we vary the debiased co-efficient $\alpha$ within the range of $\{1e\text{-}4,5e\text{-}4,1e\text{-}3,5e\text{-}3,1e\text{-}2\}$. The original retrieval models learned without the debiased constraint are denoted as “w/o debias”. The results on the SciFact+AIGC dataset are presented in Figure 8.
(a) ANCE
(b) BERM
Figure 9. Performance comparison of neural retrievers on only human-written SciFact dataset with different co-efficient $\alpha$ in our proposed debiased method.
As we can see, as the debiased co-efficient $\alpha$ increases, the Relative $\Delta$ gradually shifts from negative to positive across almost all metrics and mixed datasets. This trend indicates that the neural retrieval models can rank human-written text higher than LLM-generated text with large $\alpha$. This can be attributed to the inclusion of our debiased constraint into the learning objective, which can penalize the biased samples and compel the retrieval models not to assign higher predicted relevance scores to LLM-generated content. Moreover, as shown in Figure 9, our method not only maintains the retrieval performance on the sole human-written corpus but also provides improvements, especially with BERM as the backbone. This improvement is likely due to the inclusion of LLM-generated samples, which might enhance the model’s ability to discern relevance among similar documents.
In summary, these empirical results have demonstrated the efficacy of our proposed debiased method in mitigating source bias to different extents by adjusting the debiased coefficient $\alpha$. This flexibility allows for customizing debiasing mechanisms to meet diverse perspectives and demands. Notably, the decision to maintain equality between the two content sources or favor human-written content can be tailored based on specific requirements and environmental considerations. For example, users may not mind the content’s source if it is of high quality and fulfills their informational needs. However, bias becomes a significant issue when we aim to credit content providers and encourage more creation, impacting the sustainability of the content creation ecosystem. The optimal strategy for enhancing the sustainable development of the IR ecosystem remains an open question for further exploration.
6. Discussion: Sounding the Alarm
Through a rigorous series of experiments and thorough analysis, we have identified that neural retrieval models demonstrate clear preferences for LLM-generated texts, referred to as source bias. This bias, with the burgeoning proliferation of LLMs and AIGC, may raise significant concerns for a variety of aspects.
First, the presence of source bias poses a significant risk of gradually rendering human-written content less accessible, potentially causing a disruption in the content ecosystem. More severely, the concern is escalating with the growing prevalence of LLM-generated content online(Hanley and Durumeric, 2023; Bengio et al., 2023). Second, there is the risk that source bias may amplify the spread of misinformation, especially considering the potential of LLMs to generate deceptive content, whether intentionally or not (Chen and Shu, 2023; Su et al., 2023; Pan et al., 2023; Aslett et al., 2023). Third, source bias may be maliciously exploited to attack against neural retrieval models within today’s search engines, creating a precarious vulnerability that could be weaponized by malicious actors, reminiscent of earlier web spam link attacks against PageRank (Gyöngyi et al., 2004).
As discussed above, since LLMs can be readily instructed to generate texts at scale, source bias presents potential tangible and serious threats to the ecosystem of web content, public trust, and online safety. We hope this discussion will sound the alarm regarding the risks posed by source bias in the LLM era.
7. Related Work
Large Language Models for IR. The emergence of large language models (LLMs) (Zhao et al., 2023b; Wei et al., 2022; Yang et al., 2023) has ushered in a transformative era across various research domains, such as natural language processing (NLP) (Bubeck et al., 2023; Bang et al., 2023), education (Guo et al., 2023; Nunes et al., 2023), recommender systems (Dai et al., 2023; Fan et al., 2023), finance (Wu et al., 2023b; Huang et al., 2023), and medicine (Alberts et al., 2023; Thirunavukarasu et al., 2023). In the field of IR, much effort has also been made to utilize the remarkable knowledge and capabilities of LLMs to enhance IR systems (Ai et al., 2023; Zhu et al., 2023). In the industry community, an exemplary successful application is New Bing [^6] https://www.bing.com/new, which is an LLM-powered search assistant that adeptly extracts information from various web pages and delivers concise summarized responses to user queries, thereby improving the search experience. In the research community, there has been a proactive exploration of integrating LLMs into the IR components, including query rewriters (Srinivasan et al., 2022; Wang et al., 2023a), retrievers (Xu et al., 2022; Dai et al., 2022), re-rankers (Sun et al., 2023; Cho et al., 2023), and readers (Shi et al., 2023; Izacard et al., 2022). For a more comprehensive overview of the recent advancements in LLMs for IR, please refer to the recent survey (Zhu et al., 2023).
Artificial Intelligence Generated Content. Artificial Intelligence Generated Content (AIGC) is a rapidly advancing field that involves the creation of content using advanced Generative AI (GAI) (Cao et al., 2023; Wu et al., 2023a; Agnese et al., 2020). Unlike traditional content crafted by humans, AIGC can be generated at scale and in considerably less time (Hanley and Durumeric, 2023; Spitale et al., 2023). Recently, the development of LLMs and other GAI models has greatly improved the quality of AIGC content than before. For instance, LLMs such as ChatGPT have shown impressive abilities in generating human-like content (Cao et al., 2023; Wu et al., 2023a). The DALL-E-3 (Betker et al., 2023), another state-of-the-art text-to-image generation system, can follow user instructions to produce high-quality images. Nevertheless, as AIGC becomes more prevalent across myriad domains, ethical concerns, and potential risks come into sharper focus (Thakur et al., 2023; Wang et al., 2023b). In fact, inevitably, the GAI models may generate content with bias and discrimination as the large training data always contain bias and toxicity (Zhuo et al., 2023; Deshpande et al., 2023; Bengio et al., 2023). Furthermore, researchers have found that LLMs can be manipulated into generating increasingly deceptive misinformation, posing challenges to online safety (Jiang et al., 2023; Su et al., 2023; Chen and Shu, 2023). In addition, some recent studies indicate that training GAI models with synthetic data could result in the collapse of the next-generation models (Alemohammad et al., 2023; Shumailov et al., 2023; Briesch et al., 2023). Thus, AIGC is a double-edged sword that requires cautious handling.
8. Conclusion and Future Work
Our study offers valuable insights into several promising directions for future research, including exploring source bias in other information systems (e.g., recommender systems and advertising systems) and examining source bias in neural models towards AIGC data across multiple data modalities, not limited to text. Moreover, uncovering the root cause of the source bias and thus further mitigating it are difficult but crucial research directions.
Acknowledgements.
This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No. 62377044, 62276248, 62376275, 62076234), Beijing Natural Science Foundation (No. 4222029), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, PCC@RUC, funds for building world-class universities (disciplines) of Renmin University of China, and the Youth Innovation Promotion Association CAS under Grants No.2023111. This work was supported by the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (RUC24QSDL013). We thank all the anonymous reviewers for their positive and insightful comments.
Appendix A Example of Evaluation Metrics
In this section, we provide a toy example for illustrating the calculation of the evaluation metrics for exploring source bias, as depicted in Figure 10. Specifically, for each query, an IR model produces a ranking list that comprises documents from both human-written and LLM-generated corpora. We then calculate ranking metrics separately for human-written and LLM-generated texts, depending on the target data source. When computing metrics for one target (e.g., human-written corpus $\mathcal{C}^{H}$), the data corresponding to the other side (e.g., LLM-generated corpus $\mathcal{C}^{G}$) is disregarded (i.e., mask all the positive label $r\in\mathcal{R}^{G}$ as negative), but the rank of each document is still based on the original ranking list that incorporates a mixture of both types of text. For instance, in this toy example, when targeting the human-written corpus, the relevant document $d_{1}^{G}$ generated by LLM is treated as a negative sample. And when calculating the ranking metrics, we only consider the rank of the positive human-written documents. When the target corpus is LLM-generated, we adopt the same principle, i.e., only take the rank of the positive LLM-generated documents into account for calculating the ranking metrics.
Appendix B More Experimental Results
In Table 7, we provide the results for source bias with more common prompts sourced from InstructGPT-prompts Github Repository  [^7] https://github.com/kevinamiri/Instructgpt-prompts#rephrase-a-passage. These results indicate that common prompts can easily trigger source bias with LLM-generated content.
Table 7. Overall source bias in neural retrievers w.r.t. Relative $\Delta$ (NDCG@1) on SciFact+AIGC with mixed human-written and Llama2-generated corpora generated from different common rephrasing prompts.
Prompt ANCE BERM TAS-B Contriever
Rewrite the text below in your own words: -7.0 -22.8 -11.2 -27.6
Paraphrase the provided text while maintaining its meaning: -26.0 -55.3 -24.1 -13.6
Summarize the following passage in a concise manner: -1.4 -43.3 -34.0 -32.4
Simplify the given passage while keeping the main ideas: -29.0 -21.7 -22.5 -40.9
Rephrase the given text using alternative expressions: -25.3 -34.7 -61.9 -18.4
Condense the following passage to focus on key points: -19.0 -24.8 -22.8 -16.4
Briefly restate the provided text without losing its essence: -29.6 -50.7 -41.3 -29.8
Reword the passage below to make it more succinct: -40.6 -71.1 -54.7 -34.0
Express the following text in a different way while keeping its intent: -50.7 -39.8 -34.9 0.0
Appendix C Theoretical Analysis and Insights
@@
In the theorem below, we introduce three assumptions: Semantic Superiority, Conditional Redundancy, and Bounded Perplexity, to theoretically establish the sufficient conditions under which $\text{PPL}(d^{G},\mathcal{B})\leq\text{PPL}(d^{H},\mathcal{B})$ holds. Semantic Superiority suggests that the perplexity of human-written texts, when evaluated by humans, is lower than when evaluated by BERT. Conditional Redundancy implies that the perplexity of $d^{G}$, given $d^{H}$, is lower than the perplexity of $d^{H}$ when evaluated directly. This is intuitively true when the information added in generating from $d^{H}$ to $d^{G}$ doesn’t exceed the original information in $d^{H}$. Bounded perplexity assumes that there exists an upper bound $\epsilon$ on the increase in perplexity when evaluating $d^{G}$directly, compared to evaluating $d^{G}$ conditioned on $d^{H}$. Then we have the following theorem:
Theorem C.1.
@@
If LLM aligns more closely with BERT than with humans when predicting $d^{G}$ given $d^{H}$, such that for any $s\in[S],$
(3) $D_{\textnormal{KL}}\left(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\right)+\epsilon$
$\leq D_{\textnormal{KL}}\left(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\right),$
@@
$\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\textnormal{PPL}(d^{G},\mathcal{B})-\textnormal{PPL}(d^{H},\mathcal{B})\right]\leq 0.$
@@
In  Theorem C.1, the KL divergence is used to compare the distributions of the document $d^{G}$ conditioned on $d^{H}$ according to the LLM, BERT model, and humans. It is worth emphasizing that inequation  (LABEL:eqn:KL-distance) is not the assumption on the understanding capabilities of BERT, LLM, and humans. Instead, this inequation assumes that when predicting $d^{G}$ given $d^{H},$ the predictions by LLM are more closely aligned with those of BERT.